# Open-Vocabulary Segmentation with Unpaired Mask-Text Supervision

### v1

@@ Abstract
Contemporary cutting-edge open-vocabulary segmentation approaches commonly rely on image-mask-text triplets, yet this restricted annotation is labour-intensive and encounters scalability hurdles in complex real-world scenarios. Although some methods are proposed to reduce the annotation cost with only text supervision, the incompleteness of supervision severely limits the versatility and performance. In this paper, we liberate the strict correspondence between masks and texts by using independent image-mask and image-text pairs, which can be easily collected respectively. With this unpaired mask-text supervision, we propose a new weakly-supervised open-vocabulary segmentation framework (Uni-OVSeg) that leverages confident pairs of mask predictions and entities in text descriptions. Using the independent image-mask and image-text pairs, we predict a set of binary masks and associate them with entities by resorting to the CLIP embedding space. However, the inherent noise in the correspondence between masks and entities poses a significant challenge when obtaining reliable pairs. In light of this, we advocate using the large vision-language model (LVLM) to refine text descriptions and devise a multi-scale ensemble to stablise the matching between masks and entities. Compared to text-only weakly-supervised methods, our Uni-OVSeg achieves substantial improvements of 15.5% mIoU on the ADE20K datasets, and even surpasses fully-supervised methods on the challenging PASCAL Context-459 dataset.
Figure 1: The proposed Uni-OVSeg framework learns open-vocabulary segmentation with unpaired mask-text supervision. Compared to the labour-intensive image-mask-text annotations, independent image-mask and image-text pairs are easier to collect. With one suite of weights, given different visual prompts (e.g., points and boxes), Uni-OVSeg can segment and categorise various objects and stuff from an open set of vocabulary in the real world.
@@ 1 Introduction
Open-vocabulary segmentation refers to the segmentation and categorisation of objects from an expansive and unrestricted vocabulary, even though the object categories within the vocabulary are not encountered during training [35, 16, 70]. Compared to traditional closed-vocabulary segmentation [83, 13], which depends on predefined training categories and cannot recognise absent categories, open-vocabulary segmentation can segment any category with arbitrary text descriptions. This innovative segmentation paradigm has garnered considerable attention [67, 39] and opened up numerous potential applications [87, 88, 72].
@@
Cutting-edge approaches in open-vocabulary segmentation typically leverage supervision with triplet annotations that are composed of images, masks, and corresponding texts [63, 76]. It is worth noting that strict alignment between each mask and text results in an expensive annotation cost. To mitigate this, some weakly-supervised methods propose using only text supervision [64, 6, 65]. However, learning with this supervision, the model falls short in capturing complex spatial details, which is suboptimal for dense prediction. Furthermore, this type of supervision lacks positional information, making the model difficult to distinguish different instances with the same semantic class. These issues severely limit the versatility and segmentation performance of existing weakly-supervised methods.
@@
Motivated by that, in this paper, we propose an advanced weakly-supervised open-vocabulary segmentation framework, named Uni-OVSeg, to reduce the annotation expense while significantly enhancing performance. In essence, we liberate the strict correspondence between masks and texts, by using independent image-mask and image-text pairs. These two types of pairs can be easily collected from different sources, as illustrated in Fig. 1, such as SA-1B [34] for image-mask pairs and CC3M [50] for image-text pairs. By resorting to independent image-mask and image-text pairs, our Uni-OVSeg has a strong segmentation ability to group semantically similar pixels, and align mask-wise embeddings with entity embeddings of texts into the same space, achieving open-vocabulary segmentation.
@@
Technically, when presented with an image-mask pair, we train a visual prompt encoder, a pixel decoder, and a mask decoder to generate a set of binary masks. The collected image-text pair often contains certain texts that does not match the image [52], leading to an incorrect correspondence between masks and entities. We employ the LLaVa model [42] to refine the quality of text descriptions, coupled with a ChatGPT-based parser for precise entity extraction. For mask-text alignment, we introduce a mask-text bipartite matching to exploit confident pairs of predicted masks and entities. A multi-scale feature adapter is designed to enhance the visual embeddings of predicted masks, which are further aligned with the text embeddings of corresponding entities. As masks and text are allowed to be unpaired in Uni-OVSeg, a multi-scale ensemble is introduced to improve the quality of visual embeddings and tackle inherent noise in the correspondence between masks and entities effectively, thereby stablising the matching process. During the inference phase, a zero-shot classifier, built by embedding target dataset category names, assigns categories to predicted masks, enabling the system to segment objects across an open vocabulary. We summarise our contributions in the following aspects:
  1. 1.
@@
    We introduce a new framework Uni-OVSeg and offer an innovative solution that significantly reduces the need for costly triplet annotations. This makes open-vocabulary segmentation more accessible and scalable.
  2. 2.
    Uni-OVSeg employs a visual prompt encoder, pixel decoder, and mask decoder for binary mask segmentation. An inclusion of refinement of text descriptions, multi-scale ensemble enhances segmentation performance by alleviating inherent noise in mask-text correspondence.
  3. 3.
@@
    With one suit of weights, we achieve 32.6% mIoU on the ADE20K dataset, and even surpassing the state-of-the-art FC-CLIP [76] on the challenging PASCAL Context-459 dataset. Comprehensive ablation studies and discussions are also provided.
@@ 2 Related works
Generic segmentation. Given an image, segmentation of specific visual concepts has remained an ongoing research topic in computer vision, as indicated by the extensive literature on it [44, 23, 33]. Generic segmentation mainly includes semantic segmentation [44, 79, 77, 12], instance segmentation [23, 4, 7], and panoptic segmentation [33, 11, 56, 11], related to different levels of granularity. In more detail, semantic segmentation [20, 12, 28, 78, 38] aims to assign a label to each pixel of the input image, according to their respective semantic classes. In addition, instance segmentation [53, 59, 54] attempts to distinguish different object instances of the same semantic class. Panoptic segmentation [57, 75, 74, 9] combines the characteristics of semantic segmentation and instance segmentation. Following a close-vocabulary assumption, previous works only can predict a predefined set of object categories. In this paper, we aim to build an advanced open-vocabulary segmentation system, which can categorise objects and stuff from an open set of vocabulary in the real world.
@@
Vision foundation models. Recent advancements in visual foundation models have diversified optimisation techniques across various learning paradigms. These developments range from vision-only pretraining [24, 25, 62, 2, 61] to joint vision-language pre-training [48, 30, 73], and extend to multi-modal frameworks that integrate visual prompting [1]. A prime example of this evolution is SAM [34], which shows the potential of extensive training for general segmentation, offering impressive generalisability and scalability. Despite its impressive capabilities, SAM cannot categorise predicted masks into different semantic classes, which is limited by the supervision of the image-mask pairs. More recently, Semantic-SAM [36] unifies different sources of human-annotated segmentation datasets and augments SAM by adding semantic labels and increased levels of granularity. In our work, our aim is to develop a more flexible vision foundation model, which can be trained with unpaired mask-text supervision (e.g., independent image-mask and image-text pairs) and can be easily adapted to different segmentation tasks.
@@
Open-vocabulary segmentation. Open-vocabulary segmentation counters the constraints of closed-vocabulary segmentation by allowing the segmentation of a diverse range of classes, even those unseen during training [21, 67, 80, 68]. Existing works [17, 76, 66] leverage the pretrained vision-language models (e.g., CLIP [48] and ALIGN [30]) to perform open-vocabulary segmentation. Most open-vocabulary segmentation methods commonly utilise human-annotated supervision (i.e., the image-mask-text triplets) to generalise the capability of vision-language models from the image level to the pixel level. To reduce the dependency on this labour-intensive supervision, some weakly-supervised methods are proposed to use only text supervisions [46, 65]. They learn to group image regions into shaped segments, but struggle to distinguish different instances with the same semantic class and the segmentation performance is unsatisfactory [85, 64]. This dilemma drives our pursuit of more advanced open-vocabulary segmentation framework, where the aim is to enjoy as low annotation cost as possible and simultaneously achieve significant performance.
@@
Figure 2: Overview of the Uni-OVSeg framework. This framework consists of feature extraction, mask generation, and mask-text alignment. A frozen CLIP model and prompt encoder are used for image and text feature extraction and prompt encoding, respectively. We employs a mask and pixel decoder for binary mask prediction. A mask-text bipartite matching is designed to exploit confident mask-entity pairs. Visual prompts using boxes are omitted for simplicity.
@@ 3 Method
In this section, we first define the problem of open-vocabulary segmentation in Sec. 3.1. We then introduce a straightforward baseline in Sec. 3.2. Finally, we present our proposed Uni-OVSeg framework in Sec. 3.3, including an overview, mask generation, mask-text alignment, and open-vocabulary inference.
@@ 3.1 Problem definition
The $k$ masks $\mathbf{m}_{i}\in\{0,1\}^{H\times W}$ include the associated ground truth class $\mathbf{c}_{i}$ [76]. Unlike traditional image segmentation tasks [44, 23, 33], open-vocabulary segmentation is more challenging because inference classes are not observed during training. During the evaluation, the test categories $\mathbf{C}_{\text{test}}$ are different from $\mathbf{C}_{\text{train}}$, which contain novel categories not seen in training, i.e., $\mathbf{C}_{\text{train}}\neq\mathbf{C}_{\text{test}}$. Different from previous works [64, 66], in our setting, no paired human-annotated mask and semantic category is provided in advance for any training image. We assume that the category names of $\mathbf{C}_{\text{test}}$ are available, which are represented in the form of natural language.
@@ 3.2 Baseline
We introduce a straightforward baseline using the knowledge of image-text and image-mask pairs. Specifically, we employ a CLIP model as the visual and text encoder, which is trained on a large amount of image-text pairs. Afterward, we use the image-mask pairs to obtain a branch of mask generation, predicting a set of binary masks. To perform open-vocabulary segmentation, we crop and pool the CLIP image features based on these predicted masks, which are further classified by the CLIP text embeddings. Although this straightforward baseline enables open-vocabulary segmentation, it exhibits an obvious knowledge gap between the image-level and pixel-level tasks.
3.3 Uni-OVSeg framework
@@
Overview. An overview of our framework Uni-OVSeg, for weakly-supervised open-vocabulary segmentation is illustrated in Fig. 2. On a macro level, Uni-OVSeg contains a CLIP model to extract features of both images and text descriptions. With the image-mask pairs, a branch of mask generation, including a visual prompt encoder, a pixel decoder and a mask decoder, is employed to predict a set of binary masks of an input image. With the image-text pairs, mask-text bipartite matching is used to exploit confident pairs between predicted masks and entities in text descriptions. Afterward, we adopt a multi-scale feature adapter to enhance the mask-wise visual embeddings, which are further aligned with associated entity embeddings based on the confident pairs. Finally, we perform open-vocabulary segmentation with the above-mentioned parts. More details can be found in Appendix A.
@@
Mask generation. Given an input image $\mathbf{I}$, mask generation aims to predict a set of binary masks $\mathbf{m}$. These masks represent groups of semantically similar pixels on various locations of the image. To effectively associate these masks with visual prompts (e.g., points and boxes), we adopt an architecture inspired by [13, 34], employing a visual prompt encoder, a pixel decoder and a query-based mask decoder. Both point $(x,y)$ and box $(x,y,w,h)$ prompts are transformed into unified anchor boxes. Each point is converted into an anchor box with a predefined width $w^{{}^{\prime}}=4$ and height $h^{{}^{\prime}}=4$, effectively approximating the location of points. To address the challenge of predicting masks with varying granularities, as depicted in Fig. 2, we encode each point into two positional embeddings $\mathbf{q}^{loc}_{n}\in\mathbb{R}^{2\times dim}$ using sine encoding [55], and combine them with $M$ distinct mask embeddings $\mathbf{q}^{msk}\in\mathbb{R}^{M\times dim}$. Each mask embedding is a learnable vector corresponding to a specific granularity level. For box prompts, the process involves encoding into two position embeddings followed by concatenation with a singular mask embedding, as each box is linked to a unique object. Afterward, we represent $N$ input location prompts as a set of query embeddings $\mathbf{Q}={\mathbf{q}_{1},...,\mathbf{q}_{n}}$. For the $i$-th query,
@@
where $\mathbf{q}^{type}\in\mathbb{R}^{dim}$ denotes the query type, chosen from either the point or the box. $\mathbf{q}^{feat}\in\mathbb{R}^{dim}$ is the content embedding sampled from visual features. For box prompts, we sample visual features based on the center of each box.
Subsequently, to capture the multi-scale information, a pixel decoder equipped with multi-scale deformable attention [86] is used and outputs enhanced pixel features $\mathbf{x}^{p}$. These features, in conjunction with the query embeddings, are fed into the query-based mask decoder to predict binary masks. Each layer of the mask decoder comprises a masked cross-attention mechanism [13], a self-attention layer [55], and a feed-forward network. The binary masks for each visual prompt $\mathbf{m}_{n}$ are derived through a matrix multiplication between the query embeddings and pixel features,
@@
where $\mathbf{m}_{n}\in\mathbb{R}^{M\times H\times W}$ and $Sigmoid(\cdot)$ is the sigmoid function to normalize the mask values into $[0,1]$. An input point is likely to exist in different granularity masks simultaneously. To capture this property and accurately predict multi-granularity masks, unlike SAM [34], we devise a many-to-many matching strategy to automatically associate the ground-truth masks with the mask predictions of each point, allowing a more comprehensive use of image-mask pairs. Once matching, we compute the segmentation loss for each mask prediction $\mathbf{m}_{i}$ with the corresponding ground-truth mask $\mathbf{y}^{msk}_{j}$,
$\begin{split}\mathcal{L}_{seg}(\mathbf{m}_{i},\mathbf{y}^{msk}_{j})=&\lambda_{bce}\cdot\mathcal{L}_{bce}(\mathbf{m}_{i},\mathbf{y}^{msk}_{j})+\\ &\lambda_{dice}\cdot\mathcal{L}_{dice}(\mathbf{m}_{i},\mathbf{y}^{msk}_{j}),\end{split}$ (4)
where $\lambda_{bce}$ and $\lambda_{dice}$ are two hyper-parameters to balance the binary cross-entropy loss $\mathcal{L}_{bce}$ and the dice loss $\mathcal{L}_{dice}$. Note that we apply no penalty to the remaining unmatched mask predictions, ensuring that multiple granularity masks can be predicted for each point.
Mask-text alignment. To enable the model categorising the predicted masks from an open set of vocabulary, given the image-text pairs, we build the correspondence between objects in the image and entities in the text description. Once a set of binary masks $\mathbf{m}$ are generated from the input image, we obtain the region embeddings $\mathbf{r}_{i}$ by employing a mask pooling layer $P$ and the CLIP visual projector $F_{v}$,
$\mathbf{r}_{i}=F{v}(P(\mathbf{x}^{v},\textbf{m}_{i})),$ (5)
where $\mathbf{x}^{v}$ denotes the visual features extracted from the CLIP visual encoder. However, the collected image-text pairs often contain mismatches between visual and textual information. Directly aligning these incorrect correspondence leads to suboptimal segmentation performance. To alleviate that, we adopt a large vision-language model, LLaVa [42], to refine the text descriptions. For each text description, traditional natural language processing tools (such as NLTK) struggle to extract reliable entities due to the complex structure and diversity of natural language. Recently, ChatGPT has demonstrated strong capabilities in contextual processing, thereby more accurately identifying noun phrases. Therefore, we utilise the ChatGPT-based parser to extract reliable entities  [^1] https://openai.com/chatgpt. Afterwards, each entity combined with the prompts used in [22] is fed into the CLIP text encoder to obtain the text embedding $t_{k}$. we calculate a cost matrix $\mathbf{\Delta}$ between the region embeddings and text embeddings,
@@
where $\mathbf{\delta}^{{}^{\prime}}_{i,k}$ denotes the reverse cosine similarity between the $i$-th region embedding and the $k$-th text embedding. The scale of objects within images often varies significantly. To enhance the quality of matching, we extract region embeddings at different scales by utilising multi-resolution images. This approach allows us to integrate cost matrices from different scales to derive the final cost matrix,
$\mathbf{\Delta}^{{}^{\prime}}=\frac{1}{S}\sum_{s=1}^{S}\mathbf{\Delta}_{s},$ (7)
where $S$ denotes the number of scales and $\mathbf{\Delta}_{s}$ is the cost matrix at the $s$-th scale. With this matrix, we apply the bipartite matching algorithm [31] to obtain the confident pairs. Finally, we fed the multi-scale visual features and the CLIP visual features into a multi-scale feature adapter to obtain the advanced region embeddings $\mathbf{r}^{{}^{\prime}}_{i}$, and compute the cosine similarity loss for each paired region and text embedding,
@@
Open-vocabulary inference. During inference, given the test categories $\mathbf{C}_{\text{test}}$, we conduct prompt engineering [21] and use the CLIP text encoder to extract text embeddings for open-vocabulary segmentation. For each input image, we input an uniform point grid as the visual prompt to predict a set of binary masks, and compute the cosine similarity with the text embeddings to predict the category with the maximum similarity as the label of the corresponding mask.

### v2

@@ Abstract
Current state-of-the-art open-vocabulary segmentation methods typically rely on image-mask-text triplet annotations for supervision. However, acquiring such detailed annotations is labour-intensive and poses scalability challenges in complex real-world scenarios. While existing weakly-supervised approaches leverage image-text pairs to reduce the expansive annotation cost, the lack of mask supervision makes it difficult for the model to locate multiple instances and accurately group pixels with similar semantics, significantly hampering versatility and performance. In this paper, we introduce Unpair-Seg, a novel weakly-supervised open-vocabulary segmentation framework that learns from unpaired image-mask and image-text pairs, which can be independently and efficiently collected. Unpair-Seg initially predicts a set of binary masks and generates pseudo labels by identifying confident pairs of masks and text entities. We then train a feature adapter to align region embeddings with text embeddings based on these pseudo labels, achieving open-vocabulary segmentation. However, the inherent noise in the mask-entity correspondence poses a challenge to obtaining reliable pairs. To address this, we employ a vision-language large model to re-caption the input images and extract precise entities, and we design a multi-scale matching strategy to reduce noisy mask-entity pairs. Our Unpair-Seg framework demonstrates impressive performance, achieving 14.6% and 19.5% mIoU on the ADE-847 and PASCAL Context-459 datasets, significantly narrowing the gap between fully-supervised and weakly-supervised methods.
@@ 1 Introduction
Open-vocabulary segmentation refers to the segmentation and categorisation of objects from an expansive and unrestricted vocabulary, even though the object categories within the vocabulary are not encountered during training [33, 14]. Compared to traditional closed-vocabulary segmentation [78, 11], which depends on predefined training categories and cannot recognise absent categories, open-vocabulary segmentation can segment any category with arbitrary text descriptions. This innovative segmentation paradigm has garnered considerable attention [63, 37] and opened up numerous potential applications [82, 83, 67].
@@
Cutting-edge approaches in open-vocabulary segmentation typically leverage the supervision of triplet annotations composed of images, masks, and corresponding texts [71]. It is worth noting that strict alignment between each mask and text results in an expensive annotation cost. To mitigate this, existing weakly-supervised methods are proposed to use image-text pairs, i.e., text supervision [60, 5, 61]. However, learning with this type of supervision, the model is suboptimal to dense prediction because it is difficult to understand complex spatial relations. Furthermore, this type of supervision lacks positional information, making it challenging for the model to distinguish multiple instances with the same semantic class. The issues mentioned above severely limit these weakly-supervised methods’ versatility and segmentation performance.
Figure 1: Unpair-Seg framework directly learns from unpaired mask-text supervision. Unlike labour-intensive image-mask-text annotations, independent image-mask and image-text pairs are more accessible to collect. With a single set of weights, Unpair-Seg excels at various image segmentation tasks, including point-prompt, box-prompt, open-vocabulary semantic and panoptic segmentation. Extensive experimental results demonstrate that our method significantly narrows the gap between fully-supervised and weakly-supervised approaches.
@@
Motivated by this, we propose an advanced weakly-supervised open-vocabulary segmentation framework, Unpair-Seg, to reduce the annotation expense while significantly enhancing performance. In essence, we liberate the strict correspondence between masks and texts by using independent image-mask and image-text pairs. These two types of pairs can be easily collected from different sources, as illustrated in Figure 1, such as SA-1B [32] for image-mask pairs and LAION-400M [48] for image-text pairs. By resorting to independent image-mask and image-text pairs, our Unpair-Seg has a strong segmentation ability to group semantically similar pixels and align regional embeddings with entity embeddings of texts into the same representation space, achieving open-vocabulary segmentation.
@@
Technically, when presented with an image-mask pair, we train a visual prompt encoder, a pixel decoder, and a mask decoder to generate a set of binary masks based on different types of visual prompts (e.g., points and bounding boxes). The collected image-text pair often contains certain texts that do not match the image [51], leading to an incorrect correspondence between masks and entities. We employ the vision-language large model (i.e., InternLM-XComposer [16]) to re-caption each input image and extract more precise entities from the text description. To train mask-entity alignment, we introduce bipartite matching to exploit confident pairs of predicted masks and entities in the CLIP embedding space [46], generating pseudo labels. Due to the inherent noise in the correspondence between vision and language, we improve the quality of visual embedding by ensembling multi-scale clip features, thereby stabilising the matching process. Afterwards, given these pseudo labels, we train a feature adapter to align regional embeddings of predicted masks and entity embeddings of text descriptions. During the inference phase, a zero-shot classifier, built by embedding target dataset category names, assigns a category to each predicted mask, enabling the system to segment objects across an open vocabulary.
Before delving into details, we summarise our contributions in the following aspects:
@@
  • We introduce Unpair-Seg, a novel weakly-supervised open-vocabulary segmentation framework that significantly reduces the need for costly triplet annotations, making open-vocabulary segmentation more accessible and scalable.
  • Considering the inherent noise in the vision-language correspondence, we employ the vision-language large model to extract precise entities from input images and design a multi-scale matching strategy to exploit confident pairs of masks and entities.
@@
  • With one suit of weights, we achieve impressive results of 22.6% PQ and 19.5% mIoU on the challenging ADE20k panoptic segmentation and PASCAL Context-459 semantic segmentation, respectively. Comprehensive ablation studies and discussions are also provided.
@@ 2 Related works
Generic segmentation. Given an image, segmentation of specific visual concepts has remained an ongoing research topic in computer vision, as indicated by the extensive literature on it [42, 22, 31]. Generic segmentation mainly includes semantic segmentation [42, 74, 72, 10], instance segmentation [22, 2, 6], and panoptic segmentation [31, 9, 55, 9], related to different levels of granularity. In more detail, semantic segmentation [19, 10, 73, 36] aims to assign a label to each pixel of the input image according to their respective semantic classes. In addition, instance segmentation [52, 58, 53] attempts to distinguish different object instances of the same semantic class. Panoptic segmentation [56, 70, 69, 7] combines the characteristics of semantic segmentation and instance segmentation. Following a close-vocabulary assumption, previous works can only predict predefined object categories. In this paper, we aim to build an advanced open-vocabulary segmentation system which can categorise objects and stuff from an open set of vocabulary in the real world.
@@
Vision foundation models. Recent advancements in visual foundation models have diversified optimisation techniques across various learning paradigms. These developments range from vision-only pretraining [24, 23] to joint vision-language pre-training [46, 28, 68], and extend to multi-modal frameworks that integrate visual prompting [1]. A prime example of this evolution is SAM [32], which shows the potential of extensive training for general segmentation, offering impressive generalisability and scalability. Despite its impressive capabilities, SAM cannot categorise predicted masks into different semantic classes, which is limited by the supervision of the image-mask pairs. More recently, Semantic-SAM [34] unifies different sources of human-annotated segmentation datasets and augments SAM by adding semantic labels and increased levels of granularity. Our work aims to develop a more flexible vision foundation model, which can be trained with unpaired mask-text supervision (e.g., independent image-mask and image-text pairs) and can be easily adapted to different segmentation tasks.
@@
Open-vocabulary segmentation. Open-vocabulary segmentation counters the constraints of closed-vocabulary segmentation by allowing the segmentation of a diverse range of classes, even those unseen during training [20, 63, 75, 64]. Existing works [15, 71, 62] leverage the pretrained vision-language models (e.g., CLIP [46] and ALIGN [28]) to perform open-vocabulary segmentation. Most open-vocabulary segmentation methods commonly utilise human-annotated supervision (i.e., the image-mask-text triplets) to generalise the capability of vision-language models from the image level to the pixel level. To reduce the dependency on this labour-intensive supervision, some weakly-supervised methods are proposed to use only text supervisions [44, 61]. They learn to group image regions into shaped segments, but struggle to distinguish different instances with the same semantic class and the segmentation performance is unsatisfactory [80, 60]. This dilemma drives our pursuit of a more advanced open-vocabulary segmentation framework, where the aim is to enjoy as low an annotation cost as possible and simultaneously achieve significant performance.
@@
Figure 2: Overview of the proposed Unpair-Seg framework. Our framework consists of two stages, including mask generation and mask-entity alignment. Given image-mask pairs, we first train a prompt encoder, pixel decoder, and mask decoder for binary mask generation. Subsequently, when presented with image-text pairs, a feature adapter is optimized to align regional embeddings of predicted masks and entity embeddings of text descriptions. A mask-entity bipartite matching is designed to assign the corresponding mask prediction for each entity. CLIP visual and text encoders are frozen. Visual prompts using boxes are omitted for simplicity.
@@ 3 Method
In this section, we first define the problem of open-vocabulary segmentation in Section 3.1. We then introduce a straightforward baseline in Section 3.2. Finally, we present our proposed Unpair-Seg framework in Section 3.3, including an overview, mask generation, mask-entity alignment, and open-vocabulary inference.
@@ 3.1 Problem definition
The $k$ masks $\mathbf{m}_{i}\in\{0,1\}^{H\times W}$ include the associated ground truth class $\mathbf{c}_{i}$ [71]. Unlike traditional image segmentation tasks [42, 22, 31], open-vocabulary segmentation is more challenging because inference classes are not observed during training. During the evaluation, the test categories $\mathbf{C}_{\text{test}}$ are different from $\mathbf{C}_{\text{train}}$, which contain novel categories not seen in training, i.e., $\mathbf{C}_{\text{train}}\neq\mathbf{C}_{\text{test}}$. Different from previous works [60, 62], in our setting, no paired human-annotated mask and semantic category is provided in advance for any training image. We assume that the category names of $\mathbf{C}_{\text{test}}$ are available, represented in the form of natural language.
@@ 3.2 Baseline
We introduce a straightforward baseline using the knowledge of image-text and image-mask pairs. Specifically, we employ a CLIP model as the visual and text encoder, which is trained on a large amount of image-text pairs. Afterwards, we use the image-mask pairs to obtain a branch of mask generation, predicting a set of binary masks. To perform open-vocabulary segmentation, we crop and pool the CLIP image features based on these predicted masks, which are further classified by the CLIP text embeddings. Although this straightforward baseline enables open-vocabulary segmentation, it exhibits a noticeable knowledge gap between the image-level and pixel-level tasks.
3.3 Unpair-Seg framework
@@
Overview. Our novel framework, Unpair-Seg, for weakly-supervised open-vocabulary segmentation, is depicted in Figure 2. On a macro level, our framework contains a CLIP model to extract features of both images and text descriptions. Using the image-mask pairs, our framework employs a mask generation branch. This branch includes a visual prompt encoder, a pixel decoder, and a mask decoder, all of which work together to predict a set of binary masks for an input image. With the image-text pairs, mask-entity bipartite matching exploits confident pairs between predicted masks and entities in text descriptions. Afterwards, we adopt a multi-scale feature adapter to enhance regional embeddings of mask predictions, which are further aligned with associated entity embeddings based on the confident pairs. Finally, we perform open-vocabulary segmentation with the parts mentioned above. More details can be found in Appendix A.
@@
Mask generation. Given an input image $\mathbf{I}$, mask generation aims to predict a set of binary masks $\mathbf{m}$. These masks represent groups of semantically similar pixels at various image locations. To effectively associate these masks with visual prompts (e.g., points and boxes), we adopt an architecture inspired by [11, 32], employing a visual prompt encoder, a pixel decoder and a query-based mask decoder. Point $(x,y)$ and box $(x,y,w,h)$ prompts are transformed into unified anchor boxes. Each point is converted into an anchor box with a predefined width $w^{{}^{\prime}}$ and height $h^{{}^{\prime}}$, effectively approximating the location of points. To address the challenge of predicting masks with variable granularities, as depicted in Fig. 2, we encode each point into two positional embeddings $\mathbf{q}^{loc}_{n}\in\mathbb{R}^{2\times dim}$ using sinusoidal encoding [54], and combine them with $M$ distinct mask embeddings $\mathbf{q}^{msk}\in\mathbb{R}^{M\times dim}$. Each mask embedding is a learnable vector corresponding to a specific granularity level. For box prompts, the process involves encoding into two position embeddings followed by concatenation with a singular mask embedding, as each box is linked to a unique object. Afterward, we represent $N$ input location prompts as a set of query embeddings $\mathbf{Q}={\mathbf{q}_{1},...,\mathbf{q}_{n}}$. For the $i$-th query,
@@
where $\mathbf{q}^{type}\in\mathbb{R}^{dim}$ denotes the query type, chosen from either the point or the box. That $\mathbf{q}^{feat}\in\mathbb{R}^{dim}$ is the content embedding sampled from visual features. For box prompts, we sample visual features based on the centre of each box.
Subsequently, to capture the multi-scale information, a pixel decoder equipped with multi-scale deformable attention [81] is used and outputs enhanced pixel features $\mathbf{x}^{p}$. These features are fed into the query-based mask decoder in conjunction with the query embeddings to predict binary masks. Each layer of the mask decoder comprises a masked cross-attention layer [11], a self-attention layer [54], and a feed-forward network. The binary masks for each visual prompt $\mathbf{m}_{n}$ are derived through a matrix multiplication between the query embeddings and pixel features,
@@
where $\mathbf{m}_{n}\in\mathbb{R}^{M\times H\times W}$ and $Sigmoid(\cdot)$ is the sigmoid function to normalize the mask values into $[0,1]$. An input point is likely to exist in different granularity masks simultaneously. Given one point prompt, SAM [32] selects the prediction with the minimum cost to the ground truth for loss calculation, which cannot effectively capture the inherent ambiguity of point prompts. To model this property and accurately predict multi-granularity masks, in contrast, we design a many-to-many matching strategy to automatically associate the ground-truth masks with the mask predictions of each point, allowing a more adequate use of image-mask pairs. Once matching, we compute the segmentation loss for each mask prediction $\mathbf{m}_{i}$ with the corresponding ground-truth mask $\mathbf{y}^{msk}_{j}$. Besides, we also estimate the quality of mask generation by regressing the intersection-over-union $\mathbf{y}^{iou}_{i,j}$ between the mask prediction $\mathbf{m}_{i}$ and ground-truth mask $\mathbf{y}^{msk}_{j}$,
$\mathcal{L}_{seg}=\lambda_{bce}\cdot\mathcal{L}_{bce}(\mathbf{m}_{i},\mathbf{y}^{msk}_{j})+\lambda_{dice}\cdot\mathcal{L}_{dice}(\mathbf{m}_{i},\mathbf{y}^{msk}_{j})+\lambda_{iou}\cdot\mathcal{L}_{iou}(\mathbf{q}_{i},\mathbf{y}^{iou}_{i,j}),$ (4)
where $\lambda_{bce}$, $\lambda_{dice}$, and $\lambda_{iou}$ are three hyper-parameters to balance the binary cross-entropy loss $\mathcal{L}_{bce}$, the dice loss $\mathcal{L}_{dice}$ and the quality loss $\mathcal{L}_{iou}$.
Figure 3: Comparison between raw and improved text descriptions. “Misalign.”, “Deficient.”, and “Missing.” denote text-image misalignment, deficient description, and missing text description. Refer to caption
Note that we do not apply any penalty to the remaining unmatched predicted masks. In doing so, we ensure the model to predict multiple granularity masks of each input point prompt.
Mask-entity alignment. In order to enable the model to categorize the predicted masks from a wide range of vocabulary, we need to establish a connection between objects in the image and entities in the text description based on image-text pairs. However, these pairs often present challenges such as text-image misalignment and incomplete descriptions. Some images even lack text descriptions, making it especially difficult to align the predicted masks with the corresponding entities directly.
To address these challenges, we utilize the vision-language large model (VLLM), InternLM-XComposer [16], to re-caption each input image. Using a prompt like “You are a proficient annotator. Please output primary objects in the input image without duplication. These output objects should be connected by commas.”, we can accurately identify entities from the text descriptions, as illustrated in Figure 3.
Following this, each identified entity and the prompts used in previous work [21] are input into the CLIP text encoder to obtain the text embedding $\mathbf{t}_{k}$. Once a set of binary masks $\mathbf{m}$ is generated from the input image, we obtain the region embeddings $\mathbf{r}_{i}$ by using a mask pooling layer $P$ and the CLIP visual projector $F_{v}$. That is,
$\mathbf{r}_{i}=F_{v}(P(\mathbf{x}^{v},\textbf{m}_{i})),$ (5)
where $\mathbf{x}^{v}$ denotes the visual features extracted from the CLIP visual encoder. The scale of objects within images often varies significantly. To stabilize the process of matching mask entities across multi-resolution images, we extract regional embeddings from multi-scale dense visual features and then ensemble these regional embeddings to enhance the quality,
$\mathbf{\hat{r}}_{i}=\frac{1}{S}\sum_{s=1}^{S}\mathbf{r}_{i,s},$ (6)
where $S$ denotes the number of scales. Afterwards, we calculate a cost matrix $\mathbf{\delta}$ between the region embeddings $\mathbf{\hat{r}}$ and text embeddings $\mathbf{t}$,
@@
where $\mathbf{\delta}^{{}^{\prime}}_{i,k}$ denotes the reverse cosine similarity between the $i$-th region embedding and the $k$-th text embedding. With this matrix, we apply the bipartite matching algorithm [29] to obtain the confident pairs. Subsequently, we input the multi-scale visual features obtained from the pixel decoder and the CLIP visual features into a feature adapter to obtain the advanced region embeddings $\mathbf{r}^{{}^{\prime}}_{i}$. More details about the feature adapter can be found in Appendix A. We then compute the cosine similarity loss for each paired region and text embedding:
@@
Open-vocabulary inference. During inference, we start with the test categories $\mathbf{C}_{\text{test}}$. Then, we engage in prompt engineering [20] using the CLIP text encoder to extract text embeddings for open-vocabulary segmentation. For each input image, we use a uniform point grid as the visual prompt to predict a set of binary masks. We next calculate the cosine similarity with the text embeddings to predict the category with the highest similarity to the label of the corresponding mask.